1. If the velocity of water changes uniformly along the transition from VA = 10 m/s to VB = 4 m/s. determine the pressure difference over the distance x.



2. Air is pumped into the top of the tank so that the pressure at B is 40 kPa. Determine the discharge of water through the drainpipe when h = 5 m.


1
Expert's answer
2021-12-13T16:56:06-0500

1.

"\Delta p=\frac{\rho}2\Delta v^2=42~kPa."

2.

"Q=\frac pv=\frac p{2gh}=40~\frac ls."